OpenStudy (rleblanc19):

A vat of milk has spilled on a tile floor. The milk flow can be expressed with the function r(t) = 4t, where t represents time in minutes and r represents how far the milk is spreading. The spilled milk is creating a circular pattern on the tile. The area of the pattern can be expressed as A(r) = πr^2. Part A: Find the area of the circle of spilled milk as a function of time, or A[r(t)]. Part B: How large is the area of spilled milk after 4 minutes? You may use 3.14 to approximate π in this problem.

OpenStudy (rleblanc19):

@pooja195 @ganeshie8

OpenStudy (rleblanc19):

@KendrickLamar2014

OpenStudy (rleblanc19):

@Mehek14 @mayankdevnani

OpenStudy (rleblanc19):

@mayankdevnani @pooja195

OpenStudy (math&ing001):

For part A you plug in r(t) in A(r) A[r(t)] = π(4t)^2 = 16 π t^2 For part B you plug in 4 in A[r(t)] A[r(4)] = 16 π t^2 = 16 * 3.14 * 4^2 = ...

OpenStudy (pythagoras123):

Part A; \[r=4t\] \[a=\pi r ^{2}\] \[a=\pi (4t)^2\] \[a=16 \pi t^2\] Part B: \[a=16 \pi (4)^2\] = 803.84 units